# Artin–Schreier quandles of involutions in absolute Galois groups

### 12 Mar 2024 v1

@@ 4 Number fields
The prime field $F=\mathbb{Q}$ of rational numbers is formally real but not real closed. It has a unique ordering. In other words, the real spectrum $\operatorname{Ord}(\mathbb{Q})$ is a singleton, and all involutions in its absolute Galois group $\operatorname{Gal}(\mathbb{Q})$ form one conjugacy class. If we choose any involution $\sigma\in\mathrm{AS}(\mathbb{Q})$, then conjugation gives a homeomorphism $\operatorname{Gal}(\mathbb{Q})/\langle\sigma\rangle\cong\mathrm{AS}(\mathbb{Q})$. In particular, there are uncountably many involutions inside $\operatorname{Gal}(\mathbb{Q})$, and the topology is reasonably well understood. In contrast, the algebraic structure of the Artin–Schreier quandle $\mathrm{AS}(\mathbb{Q})$ supported on it is much more interesting:
@@
The Artin–Schreier quandle $\mathrm{AS}(\mathbb{Q})$ of the rational number field $\mathbb{Q}$ is a free pro-finite involutory quandle. A basis is given by a Cantor space of involutions inside $\operatorname{Gal}(\mathbb{Q})$.
Informally, this result says that the Artin–Schreier quandle $\mathrm{AS}(\mathbb{Q})$ is algebraically as complicated as possible, given the constraints of being an involutory quandle. Unfortunately, the proof will not display an explicit Cantor space inside $\operatorname{Gal}(\mathbb{Q})$ that is a basis for it. Instead, it implicitly uses Brouwer’s characterisation [11] of the Cantor space: every pro-finite space that meets the minimum requirements (i.e., it has to be non-empty, second countable, and does not have an isolated point) is homeomorphic to it.
@@
of the inclusion is an isomorphism of pro-finite quandles. As it is always a continuous morphism of pro-finite quandles, it will suffice to show that it is bijective.
The subset $\mathrm{AS}(\mathbb{Q})$ of involutions generates a subgroup $N$ of $\operatorname{Gal}(\mathbb{Q})$, and this subgroup is normal. Its fixed field is an infinite Galois extension $\mathbb{Q}^{\mathrm{tr}}$ of $\mathbb{Q}$, the field of totally real numbers. The normal subgroup $N$ of $\operatorname{Gal}(\mathbb{Q})$ generated by the involutions is the absolute Galois group $\operatorname{Gal}(\mathbb{Q}^{\mathrm{tr}})$ of $\mathbb{Q}^{\mathrm{tr}}$. Fried–Haran–Völklein [18, 19] and (independently) Pop [45] have shown that this absolute Galois group $\operatorname{Gal}(\mathbb{Q}^{\mathrm{tr}})$ of the field $\mathbb{Q}^{\mathrm{tr}}$ is isomorphic to the free pro-finite Coxeter group $\mathrm{F}^{\wedge}_{2}(S)$ generated by a Cantor space $S$. In other words, we can choose a subspace $S$ as above so that the canonical morphism
@@
is an isomorphism of pro-finite groups.
We will only use that (4.2) is injective. We can precompose this injection (4.2) with the canonical morphism $\mathrm{FQ}^{\wedge}_{2}(S)\to\mathrm{F}^{\wedge}_{2}(S)$, which is also injective, to obtain an injection
$\mathrm{FQ}^{\wedge}_{2}(S)\longrightarrow\mathrm{F}^{\wedge}_{2}(S)\longrightarrow\operatorname{Gal}(\mathbb{Q}^{\mathrm{tr}})\longrightarrow\operatorname{Gal}(\mathbb{Q}).$
By the universal mapping property of $\mathrm{FQ}^{\wedge}_{2}(S)$, this composition agrees with the composition
[…]
Therefore, we see that the morphism (4.1) is indeed injective.
For the proof of surjectivity of (4.1), the goal is to show that $S$ generates $\mathrm{AS}(\mathbb{Q})$ as a quandle. An arbitrary element of $\mathrm{AS}(\mathbb{Q})$ is an involution in $N\cong\mathrm{F}^{\wedge}_{2}(S)$. By Propoition 2.15, all such involutions are conjugate to generators in $S$. This fact implies that $\mathrm{AS}(\mathbb{Q})$ is generated, under conjugation, and hence as a quandle, by $S$, and the morphism (4.1) is surjective. ∎
@@
We remark that any free quandle generates a free group: the enveloping group is the value of the left-adjoint to the forgetful functor from groups to quandles, and as such, it preserves co-limits for category theoretical reasons. The proof of Theorem 4.1, however, goes the other way: it is not just a categorical formality to deduce the freeness of a quandle from the freeness of its enveloping group., and this is why any argument given necessarily has to be non-trivial.
Remark 4.3.
Our proof also used the fact that the map $\operatorname{FQ}_{2}(S)\to\mathrm{F}_{2}(S)$ is injective. This is not automatic, either. There are quandles $Q$ for which the canonical map from $Q$ into the enveloping group is not injective. For example, if $Q=\{x,y,z\}$ is the involutory quandle with $\lambda_{x}$ transposing $y$ and $z$, and $\lambda_{y}=\lambda_{z}=\operatorname{id}$, then $y\rhd x=x$ implies that $x$ commutes with $y$ in the enveloping group, and $x\rhd y=z$ implies that $x$ conjugates $y$ into $z$ in the enveloping group. Therefore, the elements $y$ and $z$ must become equal in the enveloping group.
Remark 4.4.
Passage to conjugacy classes gives a surjection
$\mathrm{AS}(F)\longrightarrow\operatorname{Ord}(F)$
of pro-finite spaces from the Artin–Schreier quandle onto the real spectrum. This surjection should not suggest, however, that we can compute the real spectrum $\operatorname{Ord}(F)$ from the Artin–Schreier quandle $\mathrm{AS}(F)$ alone. In particular, the reader may have wondered if we can recover the real spectrum of a field from its Artin–Schreier quandle as the set of orbits. (Recall that the orbits of a quandle are the equivalence classes of its elements under the equivalence relation generated by $x\rhd y\sim y$ for all $x$ and $y$.) However, this is not the case (see the following Example 4.5). Instead, we have two surjections
@@
and neither of them has to be injective. The Artin–Schreier quandle $\mathrm{AS}(F)$ is contained inside $\operatorname{Gal}(F)$, of course, and we can identify the real spectrum $\operatorname{Ord}(F)$ with the orbits of the involutions under conjugation with $\operatorname{Gal}(F)$, whereas the elements in $\operatorname{Orb}(\mathrm{AS}(F))$ are the orbits under conjugation with respect to the subgroup generated by the involutions.
@@
We have an equality $\mathrm{AS}(\mathbb{Q}(\sqrt{2}))=\mathrm{AS}(\mathbb{Q})$ of involutory quandles, as every involution in $\operatorname{Gal}(\mathbb{Q})$ also lies in the normal subgroup $\operatorname{Gal}(\mathbb{Q}(\sqrt{2}))$. However, the rational number field $\mathbb{Q}$ has a unique ordering, whereas the real quadratic extension $\mathbb{Q}(\sqrt{2})$ has two. In fact, there are involutions in $\operatorname{Gal}(\mathbb{Q}(\sqrt{2}))$ that are not conjugate in $\operatorname{Gal}(\mathbb{Q}(\sqrt{2}))$ but become conjugate in $\operatorname{Gal}(\mathbb{Q})$. The set of orbits of $\mathrm{AS}(\mathbb{Q})$ surjects onto Cantor space, whereas the real spectrum of $\mathbb{Q}$ is trivial.
Let us finally look at other examples of ‘small’ fields. The other prime fields, finite fields, and all other fields of positive characteristic are uninteresting for us as their absolute Galois groups contain no involutions. We have already discussed $\mathbb{R}$ and $\mathbb{C}$, and the other local fields of characteristic 00 (i.e., the $p$–adic fields $\mathbb{Q}_{p}$ and their finite extensions) contain no involutions either: each $\mathbb{Q}_{p}$ contains a square root of $1-p^{n}$ for sufficiently large $n$. This leaves us with the problem of extending Theorem 4.1 to other number fields.
We first establish some more notation. Let $F$ be a number field. We write $d=\dim_{\mathbb{Q}}(F)$ for its degree. Then there are isomorphisms $\mathbb{C}\otimes_{\mathbb{Q}}F\cong\mathbb{C}^{d}$ and $\mathbb{R}\otimes_{\mathbb{Q}}F\cong\mathbb{R}^{r_{1}}\oplus\mathbb{C}^{r_{2}}$ with $d=r_{1}+2r_{2}$. The $r_{1}$–element set $\operatorname{Mor}(F,\mathbb{R})$ of real embeddings $F\to\mathbb{R}$, which we can identify with its real spectrum, sits inside the $d$–element set $\operatorname{Mor}(F,\mathbb{C})$ of complex embeddings as the fixed points under the action of the Galois group $\operatorname{Aut}(\mathbb{C}|\mathbb{R})=\mathbb{Z}/2$ by complex conjugation. The orbits are the Archimedean places of $F$, with $r_{1}$ real ones and $r_{2}$ complex ones. The number field $F$ is totally real when $d=r_{1}$ and $r_{2}=0$, and $F$ is totally imaginary when $r_{1}=0$ and $d=2r_{2}$, i.e., when it is not formally real ($r_{1}>0$). If $F$ is Galois over $\mathbb{Q}$, then it has to be either of those two, as the Galois group acts transitively on the set of Archimedean places of $F$, showing that they all have to be of the same type.
@@
For any formally real number field $F$, the Artin–Schreier quandle $\mathrm{AS}(F)$ is a free pro-finite involutory quandle with a basis given by a Cantor space of involutions.
@@
If $F$ is formally real and Galois over $\mathbb{Q}$, then the group $\operatorname{Gal}(F)$ contains one involution, and it is a normal subgroup of $\operatorname{Gal}(\mathbb{Q})$. As all involutions in $\operatorname{Gal}(\mathbb{Q})$ are conjugate, we see that the group $\operatorname{Gal}(F)$ contains all involutions of $\operatorname{Gal}(\mathbb{Q})$. We deduce $\mathrm{AS}(F)=\mathrm{AS}(\mathbb{Q})$, so that in the Galois case, the claim follows trivially from Theorem 4.1.
In the general case, when $F$ is not necessarily Galois over $\mathbb{Q}$, we can proceed as in the proof of Theorem 4.1, except for one detail: we replace the field $\mathbb{Q}^{\mathrm{tr}}$ of totally real numbers and consider instead its compositum $E$ with $F$. This larger field is even better than $\mathbb{Q}^{\mathrm{tr}}$ in some sense: The field $\mathbb{Q}^{\mathrm{tr}}$ is pseudo real closed (PRC) by Pop (see [27]), but not Hilbertian. In contrast, the proper finite extensions $E$ of $\mathbb{Q}^{\mathrm{tr}}$ are PRC by Prestel [46] and Hilbertian. In fact, every proper finite extension of $\mathbb{Q}^{\mathrm{tr}}$ is Hilbertian (see [51, Satz 9.7] for a nonstandard proof and [20, Ch. 12] for a standard proof.) Hence, their absolute Galois groups
@@
are known by [21, Cor. 2]: an argument due to Jarden shows that, in the formally real case, the pro-finite group $\operatorname{Gal}(E)$ is abstractly isomorphic to $\operatorname{Gal}(\mathbb{Q}^{\mathrm{tr}})$.
The rest of the proof is nearly identical to the one for Theorem 4.1. Computing the involutions in $\operatorname{Gal}(F)$ is the same as computing the involutions in the intersection of $\operatorname{Gal}(F)$ with the subgroup generated by the involutions:
@@
From the discussion above, we know
$\operatorname{Inv}(\operatorname{Gal}(F)\cap\operatorname{Gal}(\mathbb{Q}^{\mathrm{tr}}))=\operatorname{Inv}(\operatorname{Gal}(E))\cong\operatorname{Inv}(\mathrm{F}^{\wedge}_{2}(S))$
for some Cantor space $S$, and $\operatorname{Inv}(\mathrm{F}^{\wedge}_{2}(S))=\mathrm{FQ}^{\wedge}_{2}(S)$ from Proposition 2.16. In summary, we find $\mathrm{AS}(F)\cong\mathrm{FQ}^{\wedge}_{2}(S)$, as claimed. ∎

### 08 Apr 2026 v2

@@ 4 Number fields
The prime field $F=\mathbb{Q}$ of rational numbers is formally real but not real closed. It has a unique ordering. In other words, the real spectrum $\operatorname{Ord}(\mathbb{Q})$ is a singleton, and all involutions in the absolute Galois group $\operatorname{Gal}(\mathbb{Q})$ form one conjugacy class. If we choose any involution $\sigma\in\mathrm{AS}(\mathbb{Q})$, then conjugation induces a homeomorphism $\operatorname{Gal}(\mathbb{Q})/\langle\sigma\rangle\cong\mathrm{AS}(\mathbb{Q})$, via $g\mapsto g\sigma g^{-1}$, where the subgroup $\langle\sigma\rangle$ is the centraliser of the involution $\sigma$ (see Lemma 3.5). In particular, there are uncountably many involutions inside $\operatorname{Gal}(\mathbb{Q})$, and the topology is reasonably well understood. In contrast, the algebraic structure of the Artin–Schreier quandle $\mathrm{AS}(\mathbb{Q})$ supported on it is much more interesting:
@@
The Artin–Schreier quandle $\mathrm{AS}(\mathbb{Q})$ of the rational number field $\mathbb{Q}$ is a free profinite involutory quandle. A basis is given by a Cantor space of involutions inside $\operatorname{Gal}(\mathbb{Q})$.
This result indicates that the Artin–Schreier quandle $\mathrm{AS}(\mathbb{Q})$ is algebraically as complicated as possible, given the constraints of being an involutory quandle. Unfortunately, the proof will not display an explicit Cantor space inside $\operatorname{Gal}(\mathbb{Q})$ that is a basis for it. Instead, it implicitly uses Brouwer’s characterisation [11] of the Cantor space: Every profinite space that meets the minimum requirements (i.e., it has to be non-empty, second countable, and does not have an isolated point) is homeomorphic to it.
@@
of the inclusion is an isomorphism of profinite quandles. As it is always a continuous morphism of profinite quandles, it will suffice to show that it is bijective.
The subset $\mathrm{AS}(\mathbb{Q})$ of involutions generates a normal subgroup of $\operatorname{Gal}(\mathbb{Q})$. Its closure $N$ is a closed normal subgroup with fixed field an infinite Galois extension $\mathbb{Q}^{\mathrm{tr}}$ of $\mathbb{Q}$, the field of totally real numbers, and the group $N$ can be identified with the absolute Galois group $\operatorname{Gal}(\mathbb{Q}^{\mathrm{tr}})$ of $\mathbb{Q}^{\mathrm{tr}}$. Fried–Haran–Völklein [18, 19] and (independently) Pop [50] have shown that this absolute Galois group $\operatorname{Gal}(\mathbb{Q}^{\mathrm{tr}})$ of the field $\mathbb{Q}^{\mathrm{tr}}$ is isomorphic to the free profinite Coxeter group $\mathrm{F}^{\wedge}_{2}(S)$ generated by a Cantor space $S$. In other words, we can choose a subspace $S$ as above so that the canonical morphism
@@
is an isomorphism of profinite groups. Restriction to the subset of involutions gives an isomorphism
$\operatorname{Inv}(\mathrm{F}^{\wedge}_{2}(S))\longrightarrow\operatorname{Inv}(\operatorname{Gal}(\mathbb{Q}^{\mathrm{tr}}))$
of profinite involutory quandles. The left hand side is $\mathrm{FQ}^{\wedge}_{2}(S)$ by definition (2.5). The right hand side is $\operatorname{Inv}(\operatorname{Gal}(\mathbb{Q}))$ by our choice of $\operatorname{Gal}(\mathbb{Q}^{\mathrm{tr}})$ inside $\operatorname{Gal}(\mathbb{Q})$: it contains all involutions. The result then follows from the definition $\mathrm{AS}(\mathbb{Q})=\operatorname{Inv}(\operatorname{Gal}(\mathbb{Q}))$ of the Artin–Schreier quandle. ∎
@@
For any field $F$ of characteristic zero, passage to conjugacy classes gives a surjection $\mathrm{AS}(F)\to\operatorname{Ord}(F)$ of profinite spaces from the Artin–Schreier quandle onto the real spectrum. This surjection should not suggest, however, that we can compute the real spectrum $\operatorname{Ord}(F)$ from the Artin–Schreier quandle $\mathrm{AS}(F)$ alone. In particular, the reader may have wondered if we can recover the real spectrum of a field from its Artin–Schreier quandle as the space of orbits. To make perfect sense of this question, one would first have to exercise some care to define spaces of orbits for all profinite quandles. Any reasonable definition, however, would yield the following for Artin–Schreier quandles: the group $\operatorname{Gal}(F)$ acts on the space $\mathrm{AS}(F)$ via conjugation, and $\operatorname{Orb}(\mathrm{AS}(F))$ is the space of orbits for the restriction of this action to the closed subgroup generated by $\mathrm{AS}(F)$. Yet, we still cannot recover the real spectrum of a field from its Artin–Schreier quandle as the space of orbits in this sense (see the following Example 4.3). Instead, we have two surjections.
@@
and neither of them has to be injective. The Artin–Schreier quandle $\mathrm{AS}(F)$ is contained inside $\operatorname{Gal}(F)$, of course, and we can identify the real spectrum $\operatorname{Ord}(F)$ with the orbits of the involutions under conjugation with $\operatorname{Gal}(F)$. In contrast, the elements in $\operatorname{Orb}(\mathrm{AS}(F))$ are the orbits under conjugation with respect to the closed subgroup generated by the involutions.
@@
We give an example showing that we cannot recover the real spectrum of a field from its Artin–Schreier quandle as the set of orbits. First, note that we have an equality $\mathrm{AS}(\mathbb{Q}(\sqrt{2}))=\mathrm{AS}(\mathbb{Q})$ of involutory quandles: every involution in $\operatorname{Gal}(\mathbb{Q})$ lies in the normal subgroup $\operatorname{Gal}(\mathbb{Q}(\sqrt{2}))$ (see Theorem 4.4 below for a more general statement.) Therefore, the Artin–Schreier quandles of $\mathbb{Q}(\sqrt{2})$ and $\mathbb{Q}$ are free on a Cantor space, and the space of orbits is the Cantor space in both cases. However, the rational number field $\mathbb{Q}$ has a unique ordering, whereas the real quadratic extension $\mathbb{Q}(\sqrt{2})$ has two. Some involutions in $\operatorname{Gal}(\mathbb{Q}(\sqrt{2}))$ that are not conjugate in $\operatorname{Gal}(\mathbb{Q}(\sqrt{2}))$ become conjugate in $\operatorname{Gal}(\mathbb{Q})$.
Let us finally look at other examples of ‘small’ fields. The other prime fields, finite fields, and all fields of positive characteristic are uninteresting to us, as their absolute Galois groups contain no involutions. We have already discussed $\mathbb{R}$ and $\mathbb{C}$, and the other local fields of characteristic 0 (i.e., the pp–adic fields $\mathbb{Q}_{p}$ and their finite extensions) contain no involutions either: each $\mathbb{Q}_{p}$ contains a square root of $1-p^{n}$ for sufficiently large $n$. This leaves us with the problem of extending Theorem 4.1 to other number fields.
We first establish some more notation. Let $F$ be a number field. We write $d=\dim_{\mathbb{Q}}(F)$ for its degree. Then there are isomorphisms $\mathbb{C}\otimes_{\mathbb{Q}}F\cong\mathbb{C}^{d}$ and $\mathbb{R}\otimes_{\mathbb{Q}}F\cong\mathbb{R}^{r_{1}}\oplus\mathbb{C}^{r_{2}}$ with $d=r_{1}+2r_{2}$. The $r_{1}$–element set $\operatorname{Mor}(F,\mathbb{R})$ of real embeddings $F\to\mathbb{R}$, which we can identify with its real spectrum, sits inside the dd–element set $\operatorname{Mor}(F,\mathbb{C})$ of complex embeddings as the fixed points under the action of the Galois group $\operatorname{Aut}(\mathbb{C}|\mathbb{R})=\mathbb{Z}/2$ by complex conjugation. The orbits are the Archimedean places of $F$, with $r_{1}$ real ones and $r_{2}$ complex ones. The number field $F$ is totally real when $d=r_{1}$ and $r_{2}=0$, and $F$ is totally imaginary when $r_{1}=0$ and $d=2r_{2}$. If $F$ is Galois over $\mathbb{Q}$, then it has to be either of those two, as the Galois group acts transitively on the set of Archimedean places of $F$, showing that they all have to be of the same type. However, there are totally real number fields that are not Galois over $\mathbb{Q}$.
@@
For any totally real number field $F$, the Artin–Schreier quandle $\mathrm{AS}(F)$ is a free profinite involutory quandle with a basis given by a Cantor space of involutions.
@@
In the proof of Theorem 4.1, we have used the field $\mathbb{Q}^{\mathrm{tr}}$ of totally real numbers. If the number field $F$ is totally real, then it is contained in the field $\mathbb{Q}^{\mathrm{tr}}$, and $\operatorname{Gal}(F)$ sits in between $\operatorname{Gal}(\mathbb{Q}^{\mathrm{tr}})$ and $\operatorname{Gal}(\mathbb{Q})$. In particular, these groups all have the same subsets of involutions, and we can deduce $\mathrm{AS}(F)=\mathrm{AS}(\mathbb{Q})$. ∎
In particular, if $F$ is Galois over $\mathbb{Q}$, then the Artin–Schreier quandle $\mathrm{AS}(F)$ is either empty or isomorphic to $\mathrm{AS}(\mathbb{Q})$ as described above.
One could argue that the totally real case is the most important. For instance, as Serre has remarked concerning the inverse Galois theory problem: All finite groups occur as Galois groups over $\mathbb{Q}$ if and only if all finite groups occur as Galois groups of totally real extensions of $\mathbb{Q}$ (see [32, Prop. 1]). However, there is something we can say concerning formally real number fields that are not totally real.
Proposition 4.5.
Let $F_{0}$ be a number field that is formally real, but not totally real. If $F$ is any other formally real number field that is not totally real, then we have $\mathrm{AS}(F)\cong\mathrm{AS}(F_{0})$. In words, the Artin–Schreier quandle is independent of the choice.
Proof.
We can consider the compositum $E$ of $\mathbb{Q}^{\mathrm{tr}}$ with $F$, and this field will be a proper finite extension of $\mathbb{Q}^{\mathrm{tr}}$. This larger field is even better than $\mathbb{Q}^{\mathrm{tr}}$ in some sense. While the field $\mathbb{Q}^{\mathrm{tr}}$ is pseudo real closed (PRC) by Pop (see [28]), but not Hilbertian, the proper finite extensions $E$ of $\mathbb{Q}^{\mathrm{tr}}$ are PRC by Prestel [51] and Hilbertian. In fact, every proper finite extension of $\mathbb{Q}^{\mathrm{tr}}$ is Hilbertian (see [59, Satz 9.7] for a nonstandard proof and [20, Ch. 12] for a standard proof.) Hence, their absolute Galois groups
@@
are known by [21, Cor. 2]: an argument due to Jarden shows that, in the formally real case, the profinite group $\operatorname{Gal}(E)$ is abstractly isomorphic to the free product of $\operatorname{Gal}(\mathbb{Q}^{\mathrm{tr}})\cong\mathrm{F}^{\wedge}_{2}(S)$ with a free profinite group of countable infinite rank (see [21, Cor. 1] and the remark following it). In particular, we see that the isomorphism type of $\operatorname{Gal}(E)$, while different from $\operatorname{Gal}(\mathbb{Q}^{\mathrm{tr}})$, is independent of $F$. Computing the set of involutions in $\operatorname{Gal}(F)$ is the same as computing the involutions in the intersection of $\operatorname{Gal}(F)$ with the subgroup generated by the involutions:
@@
From the discussion above, we know that $\operatorname{Inv}(\operatorname{Gal}(F)\cap\operatorname{Gal}(\mathbb{Q}^{\mathrm{tr}}))=\operatorname{Inv}(\operatorname{Gal}(E))$ is independent of $F$. ∎
Therefore, to determine the Artin–Schreier quandle of a field $F$ as above, it would be sufficient to work out one particular example. However, this is difficult for reasons related to the splitting behaviour of involutions across subgroups (see the following example). In addition, it is not known whether subquandles of free involutory quandles are free.
Example 4.6.
The absolute Galois group $\operatorname{Gal}(F)$ of a number field $F$ is a finite index subgroup of the absolute Galois group $\operatorname{Gal}(\mathbb{Q})$ of the rational number field $\mathbb{Q}$, and the reader may wonder if there is not a more abstract description of the splitting of a conjugacy class of involutions in a group $G$ when intersected with a finite index subgroup $H$ and its conjugates $gHg^{-1}$. We have not found such, but offer the following example for contemplation. Take $G$ to be the alternating group $\mathrm{A}_{5}$. It contains 15 involutions, all conjugate to $(12)(34)$. The subgroup $H$ that is generated by $(123)$ and $(12)(45)$ is isomorphic to the symmetric group $\mathrm{S}_{3}$. The group $H$ and its conjugate $gHg^{-1}$ with the element $g=(14)(25)$ both contain three involutions each, but they only have one of them in common, namely $(12)(45)$.